OpenStudy (anonymous):

3(2x + 5) ≥ 2(x + 6)

OpenStudy (anonymous):

6x+15≥2x+12 6x-2x=4x 4x+15≥12 12-15=-3 4x≥-3 -3/4=-0.75 x≥-0.75

OpenStudy (anonymous):

how about a fraction??

OpenStudy (anonymous):

-3/4

OpenStudy (anonymous):

Thank You (:

OpenStudy (anonymous):

ur welcome